I need to factor x^2+8x+16
So far i have
(x ) (x ) I know the factors of 16 are 1,2,4,8,16 and they need to add up to 8. where do i go from here?

Answer by Edwin McCravy(20054)

I need to factor x^2+8x+16
So far i have
(x ) (x ) I know the factors of 16 are 1,2,4,8,16 and they need to add up to 8. where do i go from here?

Don't just write the factors down, write them multiplied to give 16,
then add them:

16=16*1  when added give 16+1=17
16=8*2   when added give  8+2=10
16=4*4   when added give  4+4=8

So now you see that 4+4=8 is what you need, so you write:

(x+4)(x+4)

And since the factors are the same you 
write the answer in simplest form:

(x+4)²

Edwin